# Lorentzian path-integral of Robin Universe

### v1

@@ V.1 Flow equations
If one writes the complex functional as ${\cal I}=i{\cal S}/\hbar=h+iH$, and expresses $z(t)=x_{1}(t)+ix_{2}(t)$, then the downward flow can be written as
@@
where the metric $g_{ij}$ appearing above is a Riemannian metric defined on the complex plane, $\lambda$ is the flow parameter and the ()(-)( - ) sign represent the downward flow. The solution to these equations gives the steepest descent trajectories associated with the corresponding saddle point. These are also known as thimbles (denoted by ${\cal J}_{\sigma}$). If in the above set of equations the we have a (+)(+)( + ) sign instead of ()(-)( - ) then we have the equations for the steepest ascent flow lines (denoted by ${\cal K}_{\sigma}$). The subscript $\sigma$ refers to the saddle point to which these set of trajectories are attached. From this definition it is quickly seen that the real part $h$ (also called Morse function) decreases monotonically along the steepest descent lines while increasing monotonically along the steepest ascent ones. To confirm this one can consider
@@
This hold true for any generic Riemannian metric. However, for simplicity and understanding the nature of these equations, it is useful to adopt a metric (Kahler metric). In our case, we’ll opt for the flat Cartesian metric, $ds^{2}=\lvert dx\rvert^{2}$. With complex coordinates $(z,\bar{z}=(x+iy,x-iy))$, the metric components are $g_{z,z}=g_{\bar{z},\bar{z}}=0$ and $g_{z,\bar{z}}=g_{\bar{z},z}=1/2$. This simplifies flow equations as:
@@
From these set of equations one notice that the imaginary part of ${\rm Im}{\cal I}=H$ is constant along the flow lines.
@@
This feature one can exploit to find the flow trajectories which are could be hard to find out numerically as one has to deal with the differential equations. However, in order to find the nature of trajectory (whether steepest ascent/descent) one has to analyse the behaviour of ${\rm d}h(\lambda)/{\rm d}\lambda$. Note that for real ${\cal S}(z)$, the morse function $h$ will be vanish on the real line. This means that $h$ along the steepest ascent trajectories which intersect the real line will be more negative when traced backward. Hence the saddle points whose associated steepest ascent line intersect the real line will have negative $h$. Interestingly, the integration done along the thimbles of this saddle will be convergent by definition. These saddles are called relevant.
@@ V.2 Choice of contour
Once the saddle points and associated flow lines have been computed, one needs to work out way to deform the original integration contour and make the integral conditionally convergent (for more detail see Witten:2010zr ; Basar:2013eka ; Feldbrugge:2017kzv ).
@@
In the complex plane the behaviour of $h$ and $H$ will aid us in determining the correct deformed contour of integration. If one denotes a particular saddle as $N_{\sigma}$ then the region is called ‘allowed’ if for each point $N$ in the region we have $h(N)<h(N_{\sigma})$. This ‘allowed’ region is denoted by $J_{\sigma}$. A region is called ‘forbidden’ if for each point $N$ in the region we have $h(N)>h(N_{\sigma})$. The ‘forbidden’ region is denoted by $K_{\sigma}$. Intuitively, this can be thought of as landscape ridded with valleys (allowed region), hills (forbidden region) and plains (intermediate region).
It is also important to note that $h(\lambda)\to-\infty$ along the steepest descent lines when $\lambda\to\pm\infty$, while $h(\lambda)\to\infty$ along the steepest ascent lines when $\lambda\to\pm\infty$. These lines intersect only at a point which is the saddle point and where they are both well-defined
${\rm Int}\left({\cal J}_{\sigma},{\cal K}_{\sigma^{\prime}}\right)=\delta_{\sigma\sigma^{\prime}}\,.$ (83)
@@
in a homological sense for some integers $n_{\sigma}$ which will take value 00 or ±$\pm 1$ when accounting for orientation of contour over each thimble. This means $n_{\sigma}={\rm Int}({\cal C},{\cal K}_{\sigma})={\rm Int}(\mathbb{D},{\cal K}_{\sigma})$. This $n_{\sigma}$ is topological and doesn’t change if one deforms the contours. Hence the necessary and sufficient condition for a saddle to be relevant if the steepest ascent from the starting from the saddle hits the original integration contour. The corresponding thimble ${\cal J}_{\sigma}$ becomes relevant. It is also crucial that the deformed integration contour doesn’t lie in the forbidden region. In this situation it is easy to slide the original integration contour in to the deformed integration contour. Intuitively, one can think of the ease of pushing the original integration contour through the plains and valleys than crossing over the hills.
@@
It is often the case that more than one saddles are relevant and hence there will be multiple thimbles that one would need to integrate over. This will imply that the original integral has been converted in to integrals over all the relevant thimbles
$I=\int_{\cal C}{\rm d}z(t)e^{iS[z]/\hbar}=\sum_{\sigma}n_{\sigma}\int_{{\cal J}_{\sigma}}{\rm d}z(t)e^{iS[z]/\hbar}\,.$ (84)
@@
This is the feature of Picard-Lefschetz methodology. The integration done along each of the thimble is absolutely convergent as can be seen from
@@
$\biggl{|}\int_{{\cal J}_{\sigma}}{\rm d}z(t)e^{iS[z]/\hbar}\biggr{|}\leq\int_{{\cal J}_{\sigma}}\lvert{\rm d}z(t)\rvert\lvert e^{iS[z]/\hbar}\rvert=\int_{{\cal J}_{\sigma}}\lvert{\rm d}z(t)\rvert e^{h}(z)<\infty\,.$ (85)
@@
$I=\int_{\cal C}{\rm d}ze^{iS[z]/\hbar}=\sum_{\sigma}n_{\sigma}e^{iH(z_{\sigma})}\int_{{\cal J}_{\sigma}}{\rm d}ze^{h}\approx\sum_{\sigma}n_{\sigma}e^{iS[z_{\sigma}]/\hbar}\left[{\cal A}_{\sigma}+{\cal O}(\hbar)\right]\,.$ (86)
@@ VI Saddle point approximation
Having detoured a bit to learn about the Picard-Lehschetz, we return back to analyzing the mini-superspace path-integral. To proceed further and gain deeper understanding of the results obtained so far it is best to do saddle-point studies and compare it with the exact results and those obtained under the $\hbar\to 0$ limit. The $\hbar\to 0$ study has shown us that the dominant contribution in the path-integral comes from the $P_{i}=\pm 3i\sqrt{k}$, and only $P_{i}=-3i\sqrt{k}$ correspond to stable configuration. This is also the value of $P_{i}$ which leads to stable behaviour of perturbations in the Hartle-Hawking no-boundary proposal of the Universe DiTucci:2019bui ; Narain:2021bff ; Lehners:2021jmv ; DiTucci:2020weq ; DiTucci:2019dji ; Narain:2022msz . It is important to highlight that in a sense the Gauss-Bonnet favours the Hartle-Hawking no-boundary Universe.
@@ VII Connection to WDW equation
So far, we have been focussing on the path integral quantization. Employing the path integral approach is both technically and conceptually advantageous, especially when the initial condition is specified on the geometry, such as in this case, on the metric variable, rather than directly on the wave-function (as in the case of the DeWitt condition). However, one can ask how these results are related to the solution of the Wheeler-DeWitt equation in the canonical approach.
@@
where we have set $k=1$ and done integration by parts in eq. (19) to obtain the above Lagrangian. The canonical momentum $p$ corresponding to $q$ is:
@@
Consequently, the Einstein-Hilbert gravity Hamiltonian takes the form
@@
However, the Gauss-Bonnet Hamiltonian in 4 spacetime dimensions is zero Liu:2008zf i.e., ${H}_{GB}=0$., which aligns with expectations. The Gauss-Bonnet term is topological in four dimensions and should not affect the dynamics.
@@
The Wheeler-DeWitt equation corresponds to the operator version of classical constraint:
@@
where $p$, $q$ appearing in eq. (100) have been promoted to operators. The canonical commutation relation $[\hat{q},\hat{p}]=i\hbar$ and WDW can be realized in different ways.
In the following, we will employ this equation at $t=0$, where we impose the boundary condition $p_{i}+\beta q_{i}=P_{i},q_{i}=Q_{i}$. Consequently, the Hamiltonian constraint takes the form
@@
In order to derive the Wheeler-DeWitt (WDW) equation, we promote the variables to operators. It will be convenient to transition into momentum space by substituting $\hat{Q}_{i}=i\hbar\frac{\partial}{\partial P_{i}}$ and $\hat{P}_{i}=P_{i}$. Then, the WDW equation is given by
@@
where $\Psi_{0}=\Psi_{0}(P_{i})$. Note that in eq. 104, we flip the sign $\frac{\partial}{\partial P_{i}}\rightarrow-\frac{\partial}{\partial P_{i}}$ because we are considering this at the initial boundary. It is a second-order differential equation. The most general solution is given by [^3] The Airy functions $Ai$ and $Bi$ are defined as two linearly independent solutions. Rotating the arguments by $32\pi/$ 3 results in different linear combinations of Airy functions, given by: $Ai(ze^{\pm i\frac{2\pi}{3}}=\frac{1}{2}e^{\pm i\frac{\pi}{3}}\left[Ai(z)\mp iBi(z)\right]$:
@@
where, $c_{1}$ and $c_{2}$ are arbitrary functions of $\beta$ and $\Lambda$ and also $\hbar$. To match with the $t=0$ part of the path-integral expression in eq. 60, we choose the following,
@@
This fixes the constants $c_{1}$ and $c_{2}$ uniquely. At this stage, it’s worth noting a basic difference between $\beta=0$ (Neumann condition) and $\beta\neq 0$ cases. For $\beta=0$, the Wheeler-DeWitt (WDW) equation is a first-order differential equation and hence has only one solution. In this scenario, the WDW equation provides the correct wave-function without any reference to the path-integral. The reason behind this is that computing the $t=0$ part from the path-integral doesn’t require any specific choice of contour. The lesson learned is that the specific choice of contour directly relates to the arbitrary constants generated in the WDW equation. At $t=1$, it’s useful to transition to position space by substituting $p_{f}=-i\hbar\frac{\partial}{\partial q_{f}}$. Then the WDW equation reads
@@
This fixes the constants $b_{1}$, $b_{2}$ uniquely at $t=1$. Combining $t=0$ and $t=1$ parts together, we get the total wave-function $\Psi(P_{i},q_{f})=\Psi_{0}(P_{i})\Psi_{1}(q_{f})$. Hence, path integral indeed generates the solutions of the WDW equation for all values of initial momentum $P_{i}$ and $\beta$. The exponential pre-factor in eq. 60 coming from the topological Gauss-Bonnet ($\alpha$) part cannot be captured via the WDW equation. However, if we demand that the full wave function, given in eq. (60) obtained from the path-integral approach, keeping the contribution of $\alpha$ to satisfy the WDW equation, which is certainly a more stringent criterion, we notice that only for $P_{i}=-3i\sqrt{k}$ the WDW equation is satisfied at the semi-classical level. To see this, we will assume $P_{i}=-iy,\beta=-i\frac{\Lambda}{2}x$, which is the relevant range of the parameters to make the total wave-function real. Using these labellings, the WDW equation reads in the semi-classical limit:
@@
Now, assume $P_{i}\neq\pm 3i\sqrt{k}$, then eq. 110 will be satisfied at a given $(x,y)$ only for a specific value of $\alpha$, which certainly doesn’t make sense. On the other hand, for $P_{i}=\pm 3i\sqrt{k}$, the eq. 110 satisfied identically for all values of $\alpha$. From this, we conclude allowing Gauss-bonnet picks only specific values of $P_{i}$ which implement no boundary. Hence, eliminating all solutions except those no-boundary solutions. The last statement supports that no-boundary solutions give the dominant contribution in the gravitational path integral at the semi-classical level.
Now we would like to analyse the no boundary criterion (vanishing $q$) from the canonical perspective, which reads
@@
where we substituted $P_{i}=-iy$ and $\beta=-i\frac{\Lambda}{2}x$. From eq. 112, we get $y=3$ and $0\leq x\leq 1$ for the Hartle-Hawking no-boundary Universe agreeing with the previous findings. Similarly to the case of the WDW equation, demanding the entire wave-function given in eq. (60), including the Gauss-Bonnet contribution, to satisfy the above no-boundary criterion:
@@
Remarkably, again the solution $y=3$ in the interval $0\leq x\leq 1$ identically satisfies the above constraint for all values of $\alpha$, thereby realizing the Hartle-Hawking no-boundary proposal. This demonstrates that the conclusions derived from the canonical perspective are aligned with those obtained from the path integral approach, even in the presence of the Gauss-Bonnet term.

### v2

@@ V.1 Flow equations
If one writes the complex exponent as ${\cal I}=i{\cal S}/\hbar=h+iH$, and expresses $z(\lambda)=x_{1}(\lambda)+ix_{2}(\lambda)$, then the downward flow can be written as
@@
where the metric $g_{ij}$ appearing above is a Riemannian metric defined on the complex plane, $\lambda$ is the flow parameter, and the ()(-)( - ) sign represents the downward flow. The solution to these equations gives the steepest descent trajectories associated with the corresponding saddle point (also known as critical points). These are also known as thimbles (denoted by ${\cal J}_{\sigma}$, $\sigma$ subscript symbolizes saddles). If, in the above set of equations, we have a (+)(+)( + ) sign instead of ()(-)( - ), then we have the equations for the steepest ascent flow lines (denoted by ${\cal K}_{\sigma}$). The flow parameter ($\lambda$) takes the values in the range of ($\infty$- ∞,+$\infty$+ ∞). For a downward flow emanating from the saddle, $\lambda$ takes $\infty$- ∞ at the saddles and increases as we move along the flow line away from the saddles. Along the steepest ascent line, $\lambda$ increases as we move towards the saddle point and takes +$\infty$+ ∞ at the saddles. The subscript $\sigma$ refers to the saddle point to which these sets of trajectories are attached. From this definition, it is quickly seen that the real part $h$ (also called Morse function) decreases monotonically along the steepest descent lines while increasing monotonically along the steepest ascent ones. To confirm this one can consider
@@
This holds true for any generic Riemannian metric (matrices which have a positive inner product). However, for simplicity and understanding the nature of these equations, it is useful to adopt a metric (Kahler metric). In our case, we’ll opt for the flat Cartesian metric, $ds^{2}=\lvert dx\rvert^{2}$. With complex coordinates $(z,\bar{z}=(x+iy,x-iy))$, the metric components are $g_{z,z}=g_{\bar{z},\bar{z}}=0$ and $g_{z,\bar{z}}=g_{\bar{z},z}=1/2$. This simplifies flow equations as:
@@
From this set of equations, one notices that the imaginary part of ${\rm Im}{\cal I}=H$ is constant along the flow lines.
@@
This feature can be exploited to find the flow trajectories, which otherwise maybe hard to find out as one has to deal with the differential equations. However, in order to find the nature of trajectory (whether steepest ascent/descent), one just needs to analyse the behaviour of ${\rm d}h(\lambda)/{\rm d}\lambda$. Note that for real ${\cal S}(z)$, the morse function $h$ will vanish on the real line. This means that $h$ along the steepest ascent trajectories which intersect the real line will be more negative when traced backwards. Hence the saddle points whose associated steepest ascent line intersects the real line (original integration contour) will have negative $h$. Interestingly, the integration done along the thimbles of this saddle will be convergent by definition. These saddles are called relevant. In a generic situation, when $S(z)$ is not necessarily real to start with, only those saddles will be called relevant from which the associated steepest ascent lines intersect the original integration contour. These saddles contribute to the gravitational path integral and hence are called relevant.
@@ V.2 Choice of contour
Once the saddle points and associated flow lines have been computed, one needs to work out a way to deform the original integration contour and make the integral absolutely convergent (for more detail see Witten:2010zr ; Basar:2013eka ; Feldbrugge:2017kzv ).
@@
In the complex plane, the behaviour of $h$ and $H$ will aid us in determining the correct deformed contour of integration. If one denotes a particular saddle as $N_{\sigma}$ then the region is called ‘allowed’ if for each point $N$ in the region we have $h(N)<h(N_{\sigma})$. This ‘allowed’ region is denoted by $J_{\sigma}$. A region is called ‘forbidden’ if, for each point $N$ in the region, we have $h(N)>h(N_{\sigma})$. The ‘forbidden’ region is denoted by $K_{\sigma}$. Intuitively, this can be thought of as a landscape ridded with valleys (allowed region), hills (forbidden region) and plains (intermediate region).
It is also important to note that $h(\lambda)\to-\infty$ along the steepest descent lines ($\mathcal{J}_{\sigma}$) when $\lambda\to+\infty$ (i.e., moving away from the saddles), while $h(\lambda)\to\infty$ along the steepest ascent ($\mathcal{K}_{\sigma}$) lines when $\lambda\to-\infty$ (i.e., moving away from the saddles). This implies that the integrand is dying out as we move away from the saddles along the steepest descent thimbles, and the integral along the whole thimble is dominated by the saddle point contribution only. This is what we need to validate the saddle point approximation. However, along the steepest ascent lines, the integrand increases exponentially as we move away from the saddle points, and this invalidates the saddle point approximation. These lines ($\mathcal{J}_{\sigma}$ and $\mathcal{K}_{\sigma}$) intersect only at the saddle point and where they satisfy222 We have already removed the degeneracies related to the stokes ray which connects two saddle points. This happens when the steepest descent curve from a saddle point meets the steepest ascent curve from another saddle point as a result of symmetry. This can be removed by adding a small perturbation (off-shell fluctuation) to the exponent.
${\rm Int}\left({\cal J}_{\sigma},{\cal K}_{\sigma^{\prime}}\right)=\delta_{\sigma\sigma^{\prime}}\,,$ (83)
@@
where “$\rm Int$” counts the intersection between two curves. Now let us deform the original integration contour ($\mathbb{D}$) to the deformed contour ($\mathcal{C}$) and write $\mathcal{C}=\sum_{\sigma}n_{\sigma}\mathcal{J}_{\sigma}$, for some integers $n_{\sigma}$ which will take value 00 or ±$\pm 1$ depending on the orientation. This gives $n_{\sigma}={\rm Int}({\cal C},{\cal K}_{\sigma})={\rm Int}(\mathbb{D},{\cal K}_{\sigma})$. Being topological $n_{\sigma}$ doesn’t change if one deforms the contours. It also implies that the necessary and sufficient condition for a saddle to be relevant if the steepest ascent emanating from the saddle intersects the original integration contour for which only $n_{\sigma}$ is non-zero. The corresponding thimble ${\cal J}_{\sigma}$ becomes the deformed integration contour. It is also crucial that the deformed integration contour shouldn’t lie in the forbidden region near the endpoints of the original integration, which we always keep fixed during the deformation. This ensures that we can safely slide the original integration contour into the deformed integration contour by cauchy’s theorem.
@@
It is often the case that more than one saddle is relevant, and hence there will be multiple thimbles that one would need to integrate over. This will imply that the original integral has been converted into integrals over all the relevant thimbles
$I=\int_{\cal C}{\rm d}ze^{i\mathcal{S}(z)/\hbar}=\sum_{\sigma}n_{\sigma}\int_{{\cal J}_{\sigma}}{\rm d}ze^{i\mathcal{S}(z)/\hbar}\,.$ (84)
@@
This is the feature of the Picard-Lefschetz methodology. The integration done along each of the thimble is absolutely convergent as can be seen from
@@
$\biggl{|}\int_{{\cal J}_{\sigma}}{\rm d}ze^{i\mathcal{S}(z)/\hbar}\biggr{|}\leq\int_{{\cal J}_{\sigma}}\lvert{\rm d}z\rvert\lvert e^{i\mathcal{S}(z)/\hbar}\rvert=\int_{{\cal J}_{\sigma}}\lvert{\rm d}z\rvert e^{h}(z)<\infty\,.$ (85)
@@
$I=\int_{\cal C}{\rm d}ze^{i\mathcal{S}(z)/\hbar}=\sum_{\sigma}n_{\sigma}e^{iH(z_{\sigma})}\int_{{\cal J}_{\sigma}}{\rm d}ze^{h}\approx\sum_{\sigma}n_{\sigma}e^{i\mathcal{S}(z_{\sigma})/\hbar}\left[{\cal A}_{\sigma}+{\cal O}(\hbar)\right]\,.$ (86)
@@ VI Saddle point approximation
Having detoured a bit to learn about the Picard-Lehschetz, we return back to analyzing the mini-superspace path-integral. To proceed further and gain deeper understanding of the results obtained so far it is best to do saddle-point studies and compare it with the exact results and those obtained under the $\hbar\to 0$ limit. The $\hbar\to 0$ study has shown us that the dominant contribution in the path-integral comes from the $P_{i}=\pm 3i\sqrt{k}$, and only $P_{i}=-3i\sqrt{k}$ correspond to stable configuration. This is also the value of $P_{i}$, which leads to stable behaviour of perturbations in the Hartle-Hawking no-boundary proposal of the Universe DiTucci:2019bui ; Narain:2021bff ; Lehners:2021jmv ; DiTucci:2020weq ; DiTucci:2019dji ; Narain:2022msz . It is important to highlight that in a sense the Gauss-Bonnet favours the Hartle-Hawking no-boundary Universe.
@@ VII Connection to WDW equation
So far, we have been focusing on the path integral quantization. Employing the path integral approach is both technically and conceptually advantageous, especially when the initial condition is specified on the geometry, such as in this case, on the metric variable, rather than directly on the wave-function (as in the case of the DeWitt condition). However, one can ask how these results are related to the solution of the Wheeler-DeWitt equation in the canonical approach.
@@
where we have set $k=1$ and done integration by parts in eq. (19) to obtain the above Lagrangian. The canonical momentum $p$ conjugate to $q$ is:
@@
Consequently, the Hamiltonian takes the following form
@@
The Gauss-Bonnet Hamiltonian in 4 spacetime dimensions is zero Liu:2008zf i.e., ${H}_{GB}=0$. This is because the Gauss-Bonnet term is topological in four dimensions and should not affect the bulk dynamics.
@@
The Wheeler-DeWitt equation corresponds to the operator version of classical constraint $H_{EGB}=0$:
@@
where $p$, $q$ appearing in eq. (100) have been promoted to operators with the commutation relation $[\hat{q},\hat{p}]=i\hbar$. Eq. 102 implies that physical states are annihilated by the Hamiltonian.
In the following, we will employ this equation at $t=0$, where we impose the robin boundary condition $p_{i}+\beta q_{i}=P_{i},q_{i}=Q_{i}$. Consequently, the Hamiltonian constraint takes the form
@@
In order to derive the Wheeler-DeWitt (WDW) equation, we promote the variables to operators. It will be convenient to transition into momentum space by substituting $\hat{Q}_{i}=i\hbar\frac{\partial}{\partial P_{i}}$ and $\hat{P}_{i}=P_{i}$. Then, the WDW equation reads
@@
where $\Psi_{0}=\Psi_{0}(P_{i})$. Note that in eq. 104, we flip the sign $\frac{\partial}{\partial P_{i}}\rightarrow-\frac{\partial}{\partial P_{i}}$ because we are considering this at the initial boundary. It is a second-order differential equation. The most general solution is given by [^3] The Airy functions $Ai$ and $Bi$ are defined as two linearly independent solutions. Rotating the arguments by $32\pi/$ 3 results different linear combinations of Airy functions, given by: $Ai(ze^{\pm i\frac{2\pi}{3}}=\frac{1}{2}e^{\pm i\frac{\pi}{3}}\left[Ai(z)\mp iBi(z)\right]$:
@@
where, $c_{1}$ and $c_{2}$ are arbitrary functions of $\beta$ and $\Lambda$. To match with the $t=0$ part of the path-integral expression in eq. 60, we choose the following,
@@
This fixes the constants $c_{1}$ and $c_{2}$ uniquely. At this stage, it’s worth noting a basic difference between $\beta=0$ (Neumann condition) and $\beta\neq 0$ cases. For $\beta=0$, the Wheeler-DeWitt (WDW) equation is a first-order differential equation and hence has only one solution. In this scenario, the WDW equation provides the correct wave-function without any reference to the path-integral. The reason behind this is that computing the $t=0$ part from the path-integral doesn’t require any specific choice of contour. The lesson learned is that the specific choice of contour directly relates to the arbitrary constants generated in the WDW equation. At $t=1$, it’s convenient to work in position space by substituting $p_{f}=-i\hbar\frac{\partial}{\partial q_{f}}$. Then the WDW equation reads
@@
This fixes the constants $b_{1}$, $b_{2}$ uniquely at $t=1$. Combining $t=0$ and $t=1$ parts together, we get the total wave-function $\Psi(P_{i},q_{f})=\Psi_{0}(P_{i})\Psi_{1}(q_{f})$. Hence, path integral indeed generates the solutions of the WDW equation for all values of initial momentum $P_{i}$ and $\beta$ (alternatively, for a given $P_{i}$, it is a solution for all $\beta$ and vice-versa.). The exponential pre-factor in eq. 60 coming from the topological Gauss-Bonnet ($\alpha$) part cannot be captured via the WDW equation. However, if we demand that the full wave function, given in eq. (60) obtained from the path-integral approach, keeping the contribution of $\alpha$ to satisfy the WDW equation, which is certainly a more stringent criterion, we notice that only for $P_{i}=-3i\sqrt{k}$ the WDW equation is satisfied at the semi-classical level. To see this, we will assume $P_{i}=-iy,\beta=-i\frac{\Lambda}{2}x$, which is the relevant range of the parameters to make the total wave-function real. Using these relabellings, the WDW equation reads in the semi-classical limit:
@@
Now, assume $P_{i}\neq\pm 3i\sqrt{k}$, then eq. 110 will be satisfied if the term inside the square bracket vanishes. Since the $\alpha$ dependent term also contains $\Lambda$, upon vanishing, this will make $\alpha$ and $\Lambda$ dependent at arbitrarily chosen $(x,y)$ point. This certainly doesn’t make sense as $\alpha$ and $\Lambda$ are two independent parameters of our theory. Alternatively, if we take $\alpha$ and $\Lambda$ independently, then the eq. 110 equation is satisfied for a given $y$, only for specific $x$ values, which is not what we wanted. On the other hand, for $P_{i}=\pm 3i\sqrt{k}$, the eq. 110 satisfied identically for all values of $\alpha$, $\Lambda$ and $x$, which makes sense and is along the line of our expectation. From this argument, we conclude allowing Gauss-bonnet picks “only” specific values of $P_{i}$ which implement no boundary. Hence, eliminating all solutions except those no-boundary solutions. The last statement supports that no-boundary solutions give the dominant contribution to the gravitational path integral at the semi-classical level.
Now, we would like to analyse the no-boundary criterion (vanishing $q$) from the canonical perspective, which reads
@@
where we substituted $P_{i}=-iy$ and $\beta=-i\frac{\Lambda}{2}x$. From eq. 112, we get $y=3$ and $0\leq x\leq 1$ for the Hartle-Hawking no-boundary Universe agreeing with the previous findings. Similarly to the case of the WDW equation, demanding the entire wave function given in eq. (60), including the Gauss-Bonnet contribution, to satisfy the above no-boundary criterion:
@@
Remarkably, again the solution $y=3$ in the interval $0\leq x\leq 1$ identically satisfies the above constraint for all values of $\alpha$ and $\Lambda$, thereby realizing the Hartle-Hawking no-boundary proposal. This demonstrates that the conclusions derived from the canonical perspective are aligned with those obtained from the path integral approach, even in the presence of the Gauss-Bonnet term.